OpenStudy (anonymous):

Find the sum of the terms of the arithmetic sequence. 1, 7, 13, 19, . . ., 109 n = 19 A. 2204 B. 2090 C. 1045 D. 1102

hartnn (hartnn):

can u find the common difference of your sequence >?

hartnn (hartnn):

know what common difference is ?

OpenStudy (anonymous):

nope lol

hartnn (hartnn):

its the difference between current term and previous term. so what is 13-7 = ? 7-1 = ?

OpenStudy (anonymous):

13- 7 =6? 7-1=6?

hartnn (hartnn):

yes! so common difference d= 6 . now what remains is just using the formula \(\huge S_n=(n/2)(2a+(n-1)d)\) where d= 6, n= 19 is given and a =1st term =1 can u calcilate sum now ?

OpenStudy (anonymous):

im still confused

hartnn (hartnn):

confused on what? substituting the values in formula ?